Coffee cafe make 90 pounds of coffee that costs $6 per pound. The types of coffee used to make this mixture costs $7 per pound and $4 per pound. How many pounds of the $7-per-pound coffee should be used in this mixture?
Answer by mananth(16946)

Price ---------------- quantity
coffee type I $4.00 ---------------- x lb
coffee type II $7.00 ------ 90 - x lb
Mixture 6.00% ---------------- 90
Total 90 lb
4 x + 7 ( 90 - x ) = 540
4 x + 630 - 7 x = 540
4 x - 7 x = 540 - -630
-3 x = -90
/ -3
x = 30 lb $4.00 coffee type I
60 lb $7.00 coffee type II

m.ananth@hotmail.ca